For f(x) = VX+6 and g(x) = 9x + 2, find the composite function gof. (gof)(x) = Enter your answer in the answer box
Answer #1

f(x) = Nut 6, 8(x) = 94+2 gofrm=gPers) = g(+*+6) 95x46 +2 Hence ANSWER (gof) 6 = 9Jx+6 +2